Suppose you have a budgeted cost of a project at PKR 100,000,000. The project is to be completed in 12 months. After a month, you have completed 5 percent of the project at a total expense of 1,000,000. The planned completion should have been 10 percent.Using the project status calculate PV, EV, AC, BAC, CV, SPI and CPI.


1
Expert's answer
2021-02-22T17:10:25-0500

BAC 100,000,000

PV 10,000,000

EV 5,000,000

AC 1,000,000

CV 4,000,000

CPI 5

SPI 0.5


Explanation:

BAC = Budgeted cost of the project = 100,000,000

 

PV = Planned value = Planned completion percent x BAC = 10% x 10000000 = 0.10 x 100000000=10000000

 

EV = Earned value = Actual completion percent x BAC = 5% x 100,000000 = 0.05 x 100000000 = 5000000

 

AC = Actual cost incurred in the project = 1,000,000

 

CV = Cost Variance = EV - AC = 5000000 - 1000000= 4000000

 

CPI = Cost Performance Index = EV / AC = 5000000 / 1000000 = 5

 

SPI = Schedule Performance Index = EV / PV = 5000000 / 10000000 = 0.5